OpenStudy (anonymous):

vA sack holds 6 red delicious apples and 9 honey crisp apples. An apple is pulled out of the sack at random and then placed back in the sack. How many times can you expect that a red delicious apple will be pulled out of the sack in 80 tries? A. 28 B. 32 C. 40 D. 53 @misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

HI AND HELP PLEASE

OpenStudy (misty1212):

there are 15 total right? and 6 red

OpenStudy (misty1212):

compute \[\frac{6}{15}\times 80\]

OpenStudy (anonymous):

right

OpenStudy (misty1212):

or even \[\frac{2}{5}\times 80\]

OpenStudy (misty1212):

or even \[.4\times 80\] any way you like

OpenStudy (anonymous):

its 32, trust me i took the test like 5 mins ago